OpenStudy (anonymous):

Find the inverse of the func. y=log4(x+3)

OpenStudy (anonymous):

(x+3)= 4^y x= 4^y -3

OpenStudy (anonymous):

\[\log_{4}(X+3) \]

OpenStudy (anonymous):

how did you find that

OpenStudy (anonymous):

if a^x=b then loga b =x (base a)